# Self-Labeling the Job Shop Scheduling Problem

### v2

@@ Appendix E Extremely Large Instances
Table 8: Average makespan ($avg\,\pm\,std$) and execution time of CP-Sat and our SPN (when sampling $\beta=512$ solutions) on very large instances.
@@
As shown in Table 2, constructive approaches are generally inferior to state-of-the-art metaheuristics [35] (see also [50]), Mixed Integer Programming (MIP) [29], and Constraint Programming (CP) solvers. These resolution methods adopt more complex frameworks that make them more powerful for solving combinatorial problems. However, neural constructive approaches are continuously closing the gap as remarked by the SPN’s performance matching that of MIP solvers. Herein, we also prove that our SPN generally produces better results than CP solvers on very large instances.
To this end, we randomly generated 20 JSP instances with shapes $\{200\times 20,200\times 40,500\times 20,500\times 40\}$ and solved them with the CP-Sat solver of Google OR-tools 9.8, executing for 1 hour, and our SPN, sampling $\beta=512$ solutions. Table 8 reports the average execution time ($time$), the average makespan and its standard deviation ($C_{max}$) of CP-Sat and our SPN512 on each shape. The last row (Gap sum from CP-Sat) offers a relative comparison by summing up the gaps of the SPN512 computed from the solutions produced by CP-Sat. A negative value means that the SPN produced better solutions.
As one can see, the SPN512 produces better average results than CP-Sat with just a fraction of time. On these 80 instances, we observed only 7 cases in which the solutions of CP-Sat are slightly better than those of the SPN. This remarks once more the quality of our proposal, making it a valid alternative to CP solvers for very large instances.
Lastly, we remark that on these very large instances, it is possible to further reduce the execution times of our SPN by employing speed-up techniques such as model quantization (i.e., converting floats to lower precision numbers) and the new compile functionality available in PyTorch $\geq 2.0$. Our model did not undergo any of the above (herein and in any other evaluation) and has been coded by keeping the implementation as simple as possible.
@@ Appendix F Execution Times
Greedy Constructive Self-Labeling Pointer Network (SPN)
@@
Figure 5: The average execution time in seconds of the coded algorithms on different instance shapes considered in Table 2. The times of PDRs, INSA, L2D, and CL refer to the construct of a single solution. Whereas we report the times of the SPN for varying numbers of sampled solution $\beta\in\{1,32,64,128,256,512\}$.
@@
We additionally assess the timing factor, an important aspect in some scheduling scenarios. To this end, we plot in the left of Figure 5 the execution time trends of dispatching rules (PDRs), INSA, L2D (we omit CL as it has slightly higher execution times than L2D), and SPN on typical benchmark shapes, all applied in a greedy constructive manner. We also include the execution time of our SPN when sampling 512 solutions (SPN512) to prove that sampling multiple parallel solutions does not dramatically increase times. Note that all these algorithms were executed on the machine described at the beginning of Section 5. We omit other RL proposals as they do not publicly release the code, do not disclose execution times, or is unclear whether they used GPUs.
@@
When sampling 512 solutions, the SPN512 takes less than a second on medium shapes, less than 3 sec on large ones, and around 10 sec on the big $100\times 20$. As these trends align with L2D, which only constructs a single solution, our SPN is a much faster neural alternative, despite being also better in terms of quality. We also underline that the SPN and any ML-based approach are likely to be slower than PDRs and constructive heuristics, either applied in a greedy or randomized way. One can see that such approaches work faster: PDRs take 0.2 sec and INSA 1.61 sec on $100\times 20$ instances. However, this increase in execution time is not dramatic and is largely justified by better performance, especially in the case of our SPN.
@@
We finally provide in the right table of Figure 5 the precise SPN timing for varying numbers of sampled solutions $\beta$ during testing. Each row reports the average execution time on a benchmark shape and the last row (Tot Sum) sums up all the times for an overall comparison. We observe that sampling 32 and 64 solutions from the SPN takes the same except in large instances ($50\times 15,50\times 20$, and $100\times 20$ shapes), where variations are small. Therefore, sampling less than 32 solutions is pointless, as the reduction in the execution times is negligible while the quality is further reduced. Additionally, sampling more than 512 solutions further improves the SPN’s results at the cost of a consistent increment in the execution time. As an example, by sampling 1024 solutions, the overall average gap (Avg row in Table 2) is reduced by 0.4% in both benchmarks with respect to the $\beta$ = 512 case, but the execution time is roughly 1.9 times larger. Instead of blindly sampling many random solutions with long executions, it should be possible to achieve better results in less time with ad-hoc strategies, such as Beam-Search or other strategies, that sample differently based on the generated probabilities or other information. We leave this study on testing strategies to future work.
@@ Appendix G Extended Results
The extended results of the coded algorithms (PDRs, INSA, SBH, L2D, MIP, CP, and the SPN’s configurations) are available at  [^2] Extended results: https://github.com/AndreaCorsini1/SelfLabelingJobShop/blob/main/output/Results.xlsx. For computing the percentage gap (PG), we used the optimal or best-known makespan of an instance ($C_{ub}$), available at: https://optimizizer.com/jobshop.php. The extended results highlight that models trained using our Self-Labeling strategy with $\beta<256$ exhibit competitive performance and, in some instances, even surpass the model trained with $\beta=256$. This indicates the efficacy of our strategy even with a reduced number of sampled solutions. However, as remarked in the paper, sampling a larger number of training solutions yields improved overall performance, as evidenced by the lower total sums observed for larger values of $\beta$.

### v3

@@ Appendix E Extremely Large Instances
Table 8: Average makespan ($avg\,\pm\,std$) and execution time of CP-Sat and our GM (when sampling $\beta=512$ solutions) on very large instances.
@@
As shown in Tab. 2, constructive approaches are generally inferior to state-of-the-art metaheuristics [36] (see also [55]), Mixed Integer Programming (MIP) [29], and Constraint Programming (CP) solvers. These resolution methods employ more complex frameworks, making them more effective for solving combinatorial problems. However, neural constructive approaches are continuously closing the gap as remarked by the GM’s performance matching that of MIP solvers. Herein, we also prove that our GM generally produces better results than CP solvers on very large instances.
To this end, we randomly generated 20 JSP instances with shapes $\{200\times 20,200\times 40,500\times 20,500\times 40\}$ and solved them with the CP-Sat solver of Google OR-tools 9.8, executing for $1$ hour, and our GM (trained as described in Sec. 5.1), sampling $\beta=512$ solutions. Tab. 8 reports the average execution time ($time$), the average makespan and its standard deviation ($C_{max}$) of CP-Sat and our GM512 on each shape. The last row (Gap sum from CP-Sat) offers a relative comparison by summing up the gaps of the GM512 computed from the solutions produced by CP-Sat. A negative value means that the GM produced better solutions.
As one can see, the GM512 produces better average results than CP-Sat with just a fraction of time. On these $80$ instances, we observed only $7$ cases in which the solutions of CP-Sat are slightly better than those of the GM. This remarks once more the quality of our proposal, making it a valid alternative to CP solvers for extremely large instances.
Lastly, we remark that on these large instances, the execution times of our GM can be further reduced by employing speed-up techniques such as model quantization (i.e., converting floats to lower precision numbers) and the new compile functionality available in PyTorch $\geq 2.0$. Our model did not undergo any of the above (herein and in any other evaluation) and has been coded by keeping the implementation as simple as possible.
@@ Appendix F Execution Times
Greedy Constructive Our Graph Model (GM)
@@
Figure 6: The average execution time in seconds of the coded algorithms on different instance shapes considered in Tab. 2. The times of PDRs, INSA, L2D, and CL refer to the construct of a single solution. Whereas we report the times of the GM for varying numbers of sampled solution $\beta\in\{1,32,64,128,256,512\}$.
@@
We additionally assess the timing factor, an important aspect in some scheduling scenarios. To this end, we plot in the left of Fig. 6 the execution time trends of dispatching rules (PDR), INSA, L2D (we omit CL as it has slightly higher execution times than L2D), and GM on typical benchmark shapes, all applied in a greedy constructive manner. We also include the execution time of our GM when sampling 512 solutions (GM512) to prove that sampling multiple parallel solutions does not dramatically increase times. Note that all these algorithms were executed on the machine described at the beginning of Sec. 5. We omit other RL proposals as they do not publicly release the code, do not disclose execution times, or is unclear whether they used GPUs.
@@
When sampling $512$ solutions, the GM512 takes less than a second on medium shapes, less than $3$ sec on large ones, and around $10$ sec on the big $100\times 20$. As these trends align with L2D, which only constructs a single solution, our GM is a much faster neural alternative, despite being also better in terms of quality. We also underline that the GM and any ML-based approach are likely to be slower than PDRs and constructive heuristics, either applied in a greedy or randomized way. One can see that such approaches work faster: PDRs take $0.2$ sec and INSA $1.61$ sec on $100\times 20$ instances. However, this increase in execution time is not dramatic and is largely justified by better performance, especially in the case of our GM.
@@
We finally provide in the right table of Fig. 6 the precise GM timing for varying numbers of sampled solutions $\beta$ during testing. Each row reports the average execution time on a benchmark shape and the last row (Tot Sum) sums up all the times for an overall comparison. We observe that sampling $32$ and $64$ solutions from the GM takes the same except in large instances ($50\times 15,50\times 20$, and $100\times 20$ shapes), where variations are small. Therefore, sampling less than $32$ solutions is pointless, as the reduction in the execution times is negligible while the quality is further reduced. Additionally, sampling more than $512$ solutions further improves the GM’s results at the cost of a consistent increment in the execution time. As an example, by sampling $1024$ solutions, the overall average gap (Avg row in Tab. 2) is reduced by $0.4\%$ in both benchmarks with respect to the $\beta=512$ case, but the execution time is roughly $1.9$ times larger. Instead of blindly sampling many random solutions with long executions, it should be possible to achieve better results in less time with ad-hoc strategies, such as Beam-Search or other strategies, that sample differently based on the generated probabilities or other information. We leave this study on testing strategies to future work.
@@ Appendix G Extended Results
The extended results of the coded algorithms (PDRs, INSA, SBH, L2D, MIP, CP, and the GM’s configurations) are available at  [^2] Extended results: https://github.com/AndreaCorsini1/SelfLabelingJobShop/blob/main/output/Results.xlsx. For computing the percentage gap (PG), we used the optimal or best-known makespan of an instance ($C_{ub}$), available at: https://optimizizer.com/jobshop.php. The extended results highlight that models trained using our Self-Labeling strategy with $\beta<256$ exhibit competitive performance and, in some instances, even surpass the model trained with $\beta=256$. This indicates the efficacy of our strategy even with a reduced number of sampled solutions. However, as remarked in the paper, sampling a larger number of training solutions yields improved overall performance, as evidenced by the lower total sums observed for larger values of $\beta$.